1) Find the following limit:  lim (x→0) ln(1 + (sin(2x))^2 )/ (1 − cos2(x)).


2) Determine type of the differential equation y`` − 2y` + y = sin x:


◦ partial differential equation.

◦ first order differential equation.

◦ linear differential equation with constant coefficients.

◦ linear nonhomogeneous differential equation.

◦ nonlinear homogeneous differential equation.


3) Write general solution of the differential equation x^(2) y'' + xy' + a^(2) y = 0:


◦ Ax^2 + Bx

◦ Ax^a + Bx^−a

◦ Ax^ia + Bx^−ia

◦ Ae^ax + Be^−ax

◦ Ae^iax + Be^−iax

◦ explicit algebraic form does not exist.




1
Expert's answer
2021-07-19T05:48:50-0400

1)


"\lim\limits_{x\to 0}\dfrac{\ln(1+\sin(2x))^2}{1-\cos^2(x)}=\lim\limits_{x\to 0}\dfrac{2\ln(1+\sin(2x))}{\sin^2(x)}"

"\lim\limits_{x\to 0}(2\ln(1+\sin(2x))=0"

"\lim\limits_{x\to 0}(\sin^2(x))=0"

"\dfrac{0}{0}"


"\lim\limits_{x\to 0}\dfrac{\ln(1+\sin(2x))^2}{1-\cos^2(x)}=\lim\limits_{x\to 0}\dfrac{2\ln(1+\sin(2x))}{\sin^2(x)}"

"=2\lim\limits_{x\to 0}\dfrac{(\ln(1+\sin(2x)))'}{(\sin^2(x))'}"

"=2\lim\limits_{x\to 0}\dfrac{\dfrac{2\cos(2x)}{1+\sin(2x)}}{2\sin(x)\cos(x)}"

"=2\lim\limits_{x\to 0}\dfrac{\cos(2x)}{\sin(x)\cos(x)(1+\sin(2x))}"

"\lim\limits_{x\to 0^-}\dfrac{\ln(1+\sin(2x))^2}{1-\cos^2(x)}=-\infin"

"\lim\limits_{x\to 0^+}\dfrac{\ln(1+\sin(2x))^2}{1-\cos^2(x)}=+\infin"

2)

"y''− 2y' + y = \sin x"

This is linear nonhomogeneous differential equation.



3)

"x^2 y'' + xy'+ a^2 y = 0"

Let "a=0"


"x^2 y'' + xy'= 0"

"y'=z, y''=z'"


"x^2 z'+ xz= 0"

"\dfrac{dz}{z}=-\dfrac{dx}{x}"

"\ln|z|=-\ln|x|+\ln C_1"

"z=\dfrac{C_1}{x}"

"y'=\dfrac{C_1}{x}"

"y=C_1\ln x+C_2"

Therefore the answer is

explicit algebraic form does not exist.